OpenStudy (anonymous):

Given the function T(z) = 7 – z, find T(–4)

OpenStudy (lyrae):

Substitute z with -4 and evaluate.

OpenStudy (lyrae):

Example: T(-6) = 7 - (-6) = 7 + 6 = 13

OpenStudy (anonymous):

T(-4)= 11